OpenStudy (xmistermayhem):

Find the three geometric means between 2 and 162.

OpenStudy (amistre64):

do you know what a geometric sequence is?

OpenStudy (xmistermayhem):

Nope.

OpenStudy (amistre64):

then how do you usually work this sort of problem?

OpenStudy (xmistermayhem):

This is my first one. ><

OpenStudy (amistre64):

if there is a certain number of geometric means between 2 values ... say we want 2 geometric means between k and j kr^0 = k kr^1 = 1 geo mean kr^2 = 2 geo mean kr^3 = j if we know r, we can assess the geo means

OpenStudy (xmistermayhem):

Uhh... I'm sorry, I'm dumb, what do you mean?

OpenStudy (amistre64):

the list of numbers generated from k to j ... forms a geometric sequence. so 2r^4 = 162, we solve for r, to determine the 3 values that make our solution 2r^0 = 2 2r^1 2r^2 2r^3 2r^4 = 162

OpenStudy (xmistermayhem):

Oh, so it's finding what 3 numbers would go into the sequence between 2 and 162?? 2, 6, 18, 54, 162 sound right? I multipled by 3...

OpenStudy (anonymous):

let G1,G2,G3 be three G.M's between 2 and 162 so 2,G1,G2,G3,162 form a geometric sequence. no. of terms=5 first term=2 t5=162 let r be c.r. \[tn=ar ^{n-1}\] where a is first term find r then G1=ar=2r G2=G1*r G3=G2* r